OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 4)(2x + 1)(x – 7)

jimthompson5910 (jim_thompson5910):

Hint: If ABC = 0, then A = 0, B = 0 or C = 0 So if (x + 4)(2x + 1)(x – 7) = 0 then x+4=0, 2x+1=0 or x-7=0

jimthompson5910 (jim_thompson5910):

Solve each equation for x to get your 3 answers

OpenStudy (anonymous):

but im not good in algebra.

jimthompson5910 (jim_thompson5910):

x+4 = 0 x+4-4 = 0-4 x = -4 See how I got that?

OpenStudy (anonymous):

yes can you keep explaining

jimthompson5910 (jim_thompson5910):

2x+1=0 2x+1-1=0-1 2x=-1 2x/2 = -1/2 x = -1/2 So that's another zero

jimthompson5910 (jim_thompson5910):

Note: a zero is another term for root (or solution) to an equation

jimthompson5910 (jim_thompson5910):

Finally, x-7=0 x-7+7=0+7 x=7 is the third zero (or root or solution)

jimthompson5910 (jim_thompson5910):

So the three zeros are: -4, -1/2, or 7

OpenStudy (anonymous):

okay soo... x-7=0 x-7+7=0+7 x=7 rightt?

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

that's one of the zeros

OpenStudy (anonymous):

sweeeet i know how to do it now.

jimthompson5910 (jim_thompson5910):

that's great :)

OpenStudy (anonymous):

are you good in f(x)= ?